# Automatically Analyzing Performance Issues in Android Apps: How Far Are We?

### v2

@@ Abstract.
Performance issues in Android applications significantly undermine users’ experience, engagement, and retention, which is a long-lasting research topic in academia. Unlike functionality issues, performance issues are more difficult to diagnose and resolve due to their complex root causes, which often emerge only under specific conditions or payloads. Although many efforts haven attempt to mitigate the impact of performance issues by developing methods to automatically identify and resolve them, it remains unclear if this objective has been fulfilled, and the existing approaches indeed targeted on the most critical performance issues encountered in real-world settings. To this end, we conducted a large-scale comparative study of Android performance issues in real-world applications and literature. Specifically, we started by investigating real-world performance issues, their underlying root causes (i.e., contributing factors), and common code patterns. We then took an additional step to empirically summarize existing approaches and datasets through a literature review, assessing how well academic research reflects the real-world challenges faced by developers and users. Our comparison results show a substantial divergence exists in the primary performance concerns of researchers, developers, and users. Among all the identified factors, 57.14% have not been examined in academic research, while a substantial 76.39% remain unaddressed by existing tools, and 66.67% lack corresponding datasets. This stark contrast underscores a substantial gap in our understanding and management of performance issues. Consequently, it is crucial for our community to intensify efforts to bridge these gaps and achieve comprehensive detection and resolution of performance issues.
Android, Software Maintenance, Performance Issues
conference: ; ; copyright: none
@@ 1. Introduction
Our research fellows have proposed several approaches to mitigate the impact of performance issues (Li2019CharacterizingAD, ; Banerjee2014DetectingEB, ; Song2021IMGDroidDI, ; Rua2023ALE, ; Das2020CharacterizingTE, ; Vsquez2015HowDD, ; Hort2021ASO, ; Liu2019DroidLeaksAC, ; Habchi2019TheRO, ). For example, in the sense of static analysis, Li et al. (Li2019CharacterizingAD, ) developed TAPIR to analyze anti-patterns in app call graphs to detect inefficient image displays that can lead to memory consumption, energy consumption, and responsiveness issues. In addition, dynamic approaches, such as µDroid (Jabbarvand2017DroidAE, ), are proposed to detect energy waste due to resource under-utilization, wakelock misuse, and inefficient loops by monitoring abnormal energy consumption. However, given the scattered nature of current research in this field, it remains unclear whether performance issues have been comprehensively resolved. Specifically, it is unknown what the current state of Android app performance analysis is, whether current research is focused on key performance issues encountered in real-world settings, and to what extent existing methods effectively and comprehensively address these critical real-world challenges. Additionally, the capabilities of existing approaches are quite distinct, as confirmed by the fact of lacking a systematic taxonomy to map performance issues with their corresponding root causes (a.k.a, factors). Specifically, in this work, we formulate these concerns into two research questions (RQs) that we aim to answer through empirical evidence and explanatory results. The two RQs are summarized as follows.
@@
To answer these RQs, we began by conducting a large-scale analysis of real-world discussions on Android performance issues. Specifically, we collected 60,684 negative reviews from Google Play, 360,372 question threads from Stack Overflow, and 35,278 issues with 550,873 commits from GitHub to identify performance issues experienced by users and developers, analyze their root causes (i.e., contributing factors), and examine common code patterns. Following this, we conducted a systematic literature review on Android performance analysis to investigate whether existing works indeed touch on the leading performance issues encountered in real-world settings. Through a comparison of academic findings with real-world data, we developed a comprehensive taxonomy covering seven categories of performance issues and 72 contributing factors (Table 1). This taxonomy allowed us to identify substantial gaps in current academic research on Android performance analysis. Specifically, it turns out there is a substantial divergence in primary concerns among researchers, developers, and users. Additionally, it is surprising to find that among all identified factors, 57.14% have not been examined in academic research, while a substantial 76.39% remain unaddressed by existing tools, and 66.67% lack corresponding datasets. On top of these findings, we further propose future research directions to bridge these gaps.
In this paper, we make the following major contributions:
  • We conducted a large-scale exploratory study of real-world discussions on Android performance issues by analyzing crowd-sourced data from Google Play, Stack Overflow, and GitHub. This allowed us to identify the primary performance issues faced by users and developers, along with their root causes and common code patterns.
  • We conducted a systematic literature review on Android performance analysis to examine the performance issues that researchers focus on, the tools developed to address these issues, and the datasets used to evaluate the effectiveness of these tools.
  • By comparing literature with real-world data, we developed a taxonomy that categorizes seven types of performance issues and identifies 72 contributing factors, enabling us to assess how well current research objectives, tools, and datasets address real-world challenges. Based on these insights, we proposed future research directions to address the identified gaps.
@@ 2.1. Data Collection
Following common practices in analyzing large-scale online discussions (Guo2013CharacterizingAD, ; Liu2014CharacterizingAD, ; Rua2023ALE, ; Banerjee2018EnergyPatchRR, ), we collected data from three popular crowd-sourced platforms: Google Play, Stack Overflow, and GitHub. Indeed, Google Play provides user feedback to assess the impact on end-users, Stack Overflow offers discussions that reveal common issues and solutions, and GitHub contains issues and commits histories that reflect the challenges developers encounter. Fig. 1 shows our three-step pipeline to perform the real-world exploration. Briefly, we first crawl user reviews from the Google Play Store, question threads from Stack Overflow, and issue-commit pairs from GitHub. Then, we filtered this data using performance-related keywords. Lastly, we manually checked real-world performance issues, root causes, and their common code patterns.
@@ 2.1.1. Data Crawling
[Stack Overflow - Question Threads.] We also collected performance-related question threads from Stack Overflow, a widely recognized Q&A platform for programmers and developers, which plays a crucial role in the Android ecosystem by facilitating knowledge sharing, problem-solving, and community support (Wu2022RetrievingAK, ; Huang2018APIMR, ). Specifically, we explored the performance issues developers posted and discussed in the development process. Each question thread on Stack Overflow typically reflects a real-world programming challenge, with the follow-up discussions offering additional insights, perspectives, and potential solutions. To understand developers’ perspectives on Android performance issues, we crawled all question threads labeled with “android” and containing accepted answers with code, resulting in 360,372 questions.
[Github - Issues & Commits.] Additionally, we collected performance issues and commits for open-source apps from GitHub, the most widely used platform for open-source development. GitHub issues provide structured information of reported performance issues, while commits create a clear and traceable history of code fixes made as solutions. Consequently, both issues and commits offer valuable insights into how developers effectively resolve these issues during the maintenance process. To access each app’s corresponding GitHub repository, we utilized F-Droid (FDroid, ), an open-source Android app repository. We chose F-Droid because it facilitates easier inspection of code patterns in open-source applications compared to closed-source ones. To gather a representative sample of Android apps, we crawled 17 categories of open-source Android apps along with their GitHub links from F-Droid. As a result, we successfully identified 4,228 Android apps and their GitHub repositories, which contained 35,278 issues and 550,973 commits.
@@ 2.1.2. Keyword Filtering
As we focus on Android performance issues, we implemented a keyword-based filtering approach to select relevant data. To compile an accurate keyword list, we conducted a Google search for “Android performance”, reviewing the top 10 most relevant pages, which included 100 websites. Additionally, we examined several foundational papers on Android performance issues (Hort2021ASO, ; Liu2014CharacterizingAD, ; Habchi2019TheRO, ; Rua2023ALE, ; Das2020CharacterizingTE, ). From these sources, we obtained a list of 87 keywords about Android performance issues. The keyword list includes terms like “UI freeze”, “battery drain”, “memory leak”, “cpu utilization”, “app size” and so on. As a previously unknown performance issue might be caused by an existing factor and vice versa, existing keywords do not restrict us from discovering new performance issues and new factors. Using these keywords, we filtered 60,684 user reviews to 165, 360,372 Stack Overflow discussions to 893, 35,278 GitHub issues down to 317, and 550,973 GitHub commits to 780.
@@
Ultimately, we obtained 114 user reviews, 336 stack overflow discussions, 47 GitHub issues, and 295 GitHub commits. Two of the authors performed the aforementioned manual checking individually, for any disagreement, they discussed their findings until reaching a consensus. For any persisting disagreement, a senior author made the final decision. We utilized Cohen’s kappa (Landis1977AnAO, ) to assess inter-rater agreement, yielding scores of 0.944 for user reviews, 0.885 for GitHub issues, 0.895 for GitHub commits, and 0.871 for Stack Overflow posts, all of which signify substantial agreement. We then manually identified the Android performance issues within this data and present our findings in the following section.
[…]
@@ 2.2.1. Performance Issues in Real-world
Energy Consumption is also critical, as inefficient app power consumption can quickly drain batteries. These issues appeared in 21.93% of all reported reviews. E.g., one representative comment states: “severe battery drain on s10, causing 1% loss every 3 minutes. drain immediately disappeared after uninstall, back to 1% loss every 8-9 minutes.” This shows that performance issues can lead to users uninstalling the app. Storage Consumption is often complained by users that the size of the app is too large, or the cache space consumed is too large when using the app. These issues appeared in 9.65% of all reported reviews. As mobile devices have limited storage, excessive storage consumption can lead to problems like a lack of available space for other apps or files, and potential app crashes. E.g., one comment mentions: “..Every time a full size version of a Google app gets updated it eats up even more internal storage space… if there’s not enough storage space in your internal storage it won’t install certain updates or new apps from Google Play store, you have to uninstall updates or apps before reinstalling updates and apps…” Finally, Internet Data Usage issues are reported by users when apps consume excessive data, especially when they aren’t using Wi-Fi, making them more sensitive to data usage. For example, one user mentioned: “…Not sure why, but my data usage has skyrocketed. I’m on a limited plan, now I have to buy extra data every month…”
[Stack Overflow - Question Threads.] From the 336 question threads, we identified six categories of performance issues, as illustrated in the Stack Overflow pie chart in Fig. 2. Interestingly, two performance issues (i.e., CPU usage and GPU usage) are not mentioned in user reviews. Additionally, while responsiveness is the most concerning issue for end users, it’s the second most frequent topic in developer discussions. This discrepancy indicates that the performance issues developers focus on and those users are concerned about do not align. Since user satisfaction directly impacts app market share, developers should prioritize addressing the issues most important to users.
[Github - Issues & Commits.] The GitHub Issues and GitHub Commits pie charts in Fig. 2 show a similar distribution of performance issues across both data sources, with memory consumption, responsiveness, and storage consumption being the most frequently mentioned. Notably, while internet data usage is a niche issue primarily highlighted in user reviews and absent from Stack Overflow and GitHub commits, it also appears in GitHub issues, providing insights into its underlying causes.
@@
Table 1. Taxonomy of seven categories of performance issues and their contributing factors. ‘R-W.’ indicates real-world sources, while ‘Lit.’ refers to literature sources.
Category Factor R-W. Lit. Problem Example
@@
Image Decoding images without resizing; leaking image objects  (Song2021IMGDroidDI, ; Li2019CharacterizingAD, ; Carette2017InvestigatingTE, ; Kang2016ExperienceRD, )
@@
Resource Improper releasing of resources (e.g., Sensor, Camera)  (Liu2014GreenDroidAD, ; Liu2013WhereHM, ; Li2019DetectingAD, ; Behrouz2016EnergyawareTM, ; Li2022CombattingEI, ; Liu2016FixingRL, ; Couto2020EnergyRF, ; Jabbarvand2020AutomatedCO, ; Li2020EnergyDxDE, ; Lu2022DetectingRU, ; Banerjee2018EnergyPatchRR, ; Wu2016LightWeightIA, ; Jabbarvand2019SearchBasedET, ; Banerjee2014DetectingEB, ; Yan2013SystematicTF, ; Liu2019DroidLeaksAC, )
@@
View Making allocations inside onDraw() routines  (Habchi2019TheRO, ; Couto2020EnergyRF, )
@@
Data Type Declaring a variable as static holds a strong reference to the previous task, preventing garbage collection (Das2020CharacterizingTE, )
@@
Inefficient Method Selecting a less efficient method introduces unnecessary delays in processing UI actions
@@
Resource Insufficient thread resource allocation causes slow response during simultaneous operations (Bhatt2020AutomatedRO, )
Image 1) Decoding large images in the UI thread, 2) Loading large image, 3) dynamically adjusting content during activity-switching events (Song2021IMGDroidDI, ; Li2019CharacterizingAD, ; Carette2017InvestigatingTE, ; Kang2016ExperienceRD, )
@@
UI Rendering Improperly stacking and layering in layouts can result in pixels being redrawn multiple times in the same frame, with only the last rendering visible (Gao2017EveryPC, ; Lin2024AgingOG, ; Liu2014CharacterizingAD, ; Habchi2019TheRO, ; Vsquez2015HowDD, ; Rua2023ALE, ; Das2020CharacterizingTE, ; Janssen2022OnTI, )
Loop An infinite while loop in the button’s onClick() event makes the app unresponsive after clicking (Lyu2018RemoveRF, )
Main Thread Performing heavy operations in the main thread  (Liu2014CharacterizingAD, ; Brocanelli2018HangDR, ; Lin2014RetrofittingCF, ; Gmez2016MiningTR, ; Vsquez2015HowDD, ; Das2020CharacterizingTE, )
Activity Loading new resources during activity restarts may freeze the UI  (Guo2022DetectingAF, ; Chen2023TransparentRC, ; Lu2022DetectingRU, )
AsyncTask Processing tasks in a single-threaded pool of AsyncTask since Android 3.0 (Lin2015StudyAR, ; Pan2020StaticAC, ; Kang2016DiagDroidAP, ; Fan2018EfficientlyMA, )
View Allocating new objects along with draw operations (Habchi2019TheRO, ; Couto2020EnergyRF, ; Das2020CharacterizingTE, )
@@
Message Backlogging rapidly incoming messages in Message Handler’s single-threaded (Kang2016DiagDroidAP, )
Storage Consumption File Handling Temp files from URIs aren’t cleared after upload

### v3

# A Comparative Study of Android Performance Issues in Real-world Applications and Literature

@@ Abstract.
Performance issues in Android applications significantly undermine users’ experience, engagement, and retention, which is a long-lasting research topic in academia. Unlike functionality issues, performance issues are more difficult to diagnose and resolve due to their complex root causes, which often emerge only under specific conditions or payloads. Although many efforts have attempted to mitigate the impact of performance issues by developing methods to automatically identify and resolve them, it remains unclear if this objective has been fulfilled, and the existing approaches indeed targeted on the most critical performance issues encountered in real-world settings. To this end, we conduct a large-scale comparative study of Android performance issues in real-world applications and literature. Specifically, we started by investigating real-world performance issues, their underlying root causes (i.e., contributing factors), and common code patterns. We then took an additional step to empirically summarize existing approaches and datasets through a literature review, assessing how well academic research reflects the real-world challenges faced by developers and users. Our comparison results show a substantial divergence exists in the primary performance concerns of researchers, developers, and users. Among all the identified factors, 57.14% have not been examined in academic research, while a substantial 63.41% remain unaddressed by existing tools, and 70.73% lack corresponding datasets. This stark contrast underscores a substantial gap in our understanding and management of performance issues. Consequently, it is crucial for our community to intensify efforts to bridge these gaps and achieve comprehensive detection and resolution of performance issues.
Android, Software Maintenance, Performance Issues, Comparative Study
$c$s: Software and its engineering
@@ 1. Introduction
To address performance issues in Android applications, researchers have explored the topic from multiple perspectives. On one hand, several studies aim to understand real-world performance issues by analyzing practical data sources (Kumari2024AnES, ; Noor2022EndUP, ; Das2016AQA, ; Vsquez2015HowDD, ). For instance, Noor et al. (Noor2022EndUP, ) examined 368,704 low-rated Google Play reviews and identified eight categories of user-reported performance issues, such as responsiveness delays and graphical glitches. Kumari et al. (Kumari2024AnES, ) analyzed 385 Stack Overflow posts to summarize developer-discussed performance issues and their causes, including UI/UX inefficiencies and problematic development practices. On the other hand, academic research has focused on performance-related code smells and design flaws (LeGoaer2022ecoCodeAS, ; Wu2023ASL, ; Fawad2024AndroidSC, ; Palomba2019OnTI, ). For example, Wu et al. (Wu2023ASL, ) reviewed 35 studies and categorized Android code smells into five types, including energy inefficiency and memory overuse. Similarly, Fawad et al. (Fawad2024AndroidSC, ) examined 79 papers and identified 237 Android code smell types, many of which are closely related to performance concerns, such as wakelock misuse and inefficient database queries. In addition to identifying performance issues, researchers have proposed various techniques to mitigate them (Li2019CharacterizingAD, ; Banerjee2014DetectingEB, ; Song2021IMGDroidDI, ; Rua2023ALE, ; Das2020CharacterizingTE, ; Vsquez2015HowDD, ; Hort2021ASO, ; Liu2019DroidLeaksAC, ; Habchi2019TheRO, ). For example, in the area of static analysis, Li et al. (Li2019CharacterizingAD, ) developed TAPIR, which analyzes anti-patterns in app call graphs to detect inefficient image displays that can lead to memory consumption, energy consumption, and responsiveness issues. Meanwhile, dynamic approaches such as µDroid (Jabbarvand2017DroidAE, ) have been proposed to detect energy waste caused by resource underutilization, wakelock misuse, and inefficient loops by monitoring abnormal energy consumption.
However, despite extensive research efforts, the study of Android performance issues remains fragmented. Most existing studies focus on a single data source or perspective, such as the user perspective from Google Play (Noor2022EndUP, ), the developer perspective from GitHub issues (Das2016AQA, ; Vsquez2015HowDD, ; Liu2014CharacterizingAD, ), or the researcher perspective from academic literature (Wu2023ASL, ; Fawad2024AndroidSC, ; Hort2021ASO, ). Moreover, these studies examine performance issues from different dimensions: some focus on symptoms (e.g., memory leaks (Banerjee2018EnergyPatchRR, ; Liu2016FixingRL, )), others on root causes (e.g., improper wakelock management (Pathak2012WhatIK, )), or specific code-level patterns (e.g., inefficient image loading (Song2021IMGDroidDI, )). Terminological inconsistencies further exacerbate this fragmentation, as similar issues are defined differently across studies, such as green bugs (Goar2020EnforcingGC, ), energy leaks (Carette2017InvestigatingTE, ), and energy smells (Iannone2020RefactoringAE, ). As a result, it remains unclear what the current state of Android app performance analysis is, whether current research is focused on key performance issues encountered in real-world settings, and to what extent existing methods effectively and comprehensively address these critical real-world challenges. Specifically, in this work, we formulate these concerns into two research questions (RQs) that we aim to answer through empirical evidence and explanatory results. The two RQs are summarized as follows.
@@
To answer these RQs, we began by conducting a large-scale analysis of real-world discussions on Android performance issues. Specifically, we collected 60,684 negative reviews from Google Play, 749,067 question threads from Stack Overflow, and 16,977 issues with 344,922 commits from GitHub to identify performance issues experienced by users and developers, analyze their root causes (i.e., contributing factors), and examine common code patterns. Following this, we conducted a systematic literature review on Android performance analysis to investigate whether existing works indeed touch on the leading performance issues encountered in real-world settings. Through a comparison of academic findings with real-world data, we developed a comprehensive taxonomy covering seven categories of performance issues and 82 contributing factors (Table 1). This taxonomy (1) unifies inconsistent terminologies that fragmented prior studies, (2) explicitly distinguishes consequences from causes, providing a clearer mapping between real-world problems and their contributing factors, and (3) integrates user, developer, and researcher perspectives within a single framework. Together, these features enable systematic cross-perspective analysis and allow us to identify substantial gaps in current academic research on Android performance analysis. Specifically, it turns out there is a substantial divergence in primary concerns among researchers, developers, and users. Additionally, it is surprising to find that among all identified factors, 57.14% have not been examined in academic research, while a substantial 63.41% remain unaddressed by existing tools, and 70.73% lack corresponding datasets. On top of these findings, we further propose future research directions to bridge these gaps.
In this paper, we made the following key contributions:
  • Cross-perspective analysis of real-world and academic sources. We collected and analyzed performance-related data from Google Play (user perspective), Stack Overflow and GitHub (developer perspective), and academic literature (researcher perspective), providing the first unified investigation across real-world and academic sources from multiple stakeholder perspectives.
  • A unified consequence–cause taxonomy. We proposed a hierarchical taxonomy that systematically links observable performance issues (e.g., sluggish UI, memory bloat) to their contributing factors (e.g., inefficient image handling, unreleased wakelocks). The taxonomy covers seven types of performance issues and identifies 82 contributing factors, enabling consistent classification and comparison across sources.
  • Quantitative gap analysis between research and practice. Using the taxonomy, we compared literature with real-world data to assess how well current research, tools, and datasets align with real-world concerns, and proposed directions to address the identified gaps.
  • Artifact release for reproducibility. We have released our artifacts (our_repo, ), including source code, dataset, and summarized results, to support future research and community development.
@@ 2.1. Data Collection
Following common practices in analyzing large-scale online discussions (Guo2013CharacterizingAD, ; Liu2014CharacterizingAD, ; Rua2023ALE, ; Banerjee2018EnergyPatchRR, ), we collected data from three popular crowd-sourced platforms: Google Play, Stack Overflow, and GitHub. Indeed, Google Play provides user feedback to assess the impact on end-users, Stack Overflow offers discussions that reveal common issues and solutions, and GitHub contains issues and commit histories that reflect the challenges developers encounter. Fig. 1 shows our three-step pipeline to perform the real-world exploration. Briefly, we first crawl user reviews from the Google Play Store, question threads from Stack Overflow, and issue-commit pairs from GitHub. Then, we filtered this data using performance-related keywords. Lastly, we manually checked real-world performance issues, root causes, and their common code patterns.
@@ 2.1.1. Data Crawling
[Stack Overflow - Question Threads.] We also collected performance-related question threads from Stack Overflow, a widely recognized Q&A platform for programmers and developers, which plays a crucial role in the Android ecosystem by facilitating knowledge sharing, problem-solving, and community support (Wu2022RetrievingAK, ; Huang2018APIMR, ). Specifically, we explored the performance issues developers posted and discussed in the development process. Each question thread on Stack Overflow typically reflects a real-world programming challenge, with the follow-up discussions offering additional insights, perspectives, and potential solutions. To understand developers’ perspectives on Android performance issues, we crawled all question threads labeled with “android” with code, resulting in 749,067 questions.
[Github - Issues & Commits.] Additionally, we collected performance issues and commits for open-source apps from GitHub, the most widely used platform for open-source development. GitHub issues provide structured information of reported performance issues, while commits create a clear and traceable history of code fixes made as solutions. Consequently, both issues and commits offer valuable insights into how developers effectively resolve these issues during the maintenance process. To access each app’s corresponding GitHub repository, we utilized F-Droid (FDroid, ), an open-source Android app repository. We chose F-Droid because it facilitates easier inspection of code patterns in open-source applications compared to closed-source ones. To gather a representative sample of Android apps, we crawled 17 categories of open-source Android apps along with their GitHub links from F-Droid, resulting in an initial dataset of 4,228 open-source Android apps. To ensure that the selected apps are not dummy projects, we introduced an additional validation step. Specifically, we verified whether each app also had a corresponding link to the Google Play Store. This yielded 1,643 apps that were both open-source and publicly distributed. From the repositories of these 1,643 apps, we extracted a total of 16,977 GitHub issues and 344,922 commits for our subsequent analysis.
@@ 2.1.2. Keyword Filtering
As we focus on Android performance issues, we implemented a keyword-based filtering approach to select relevant data. To compile an accurate keyword list, we conducted a Google search for “Android performance”, reviewing the top 10 most relevant pages, which included 100 websites. Additionally, we examined several foundational papers on Android performance issues (Hort2021ASO, ; Liu2014CharacterizingAD, ; Habchi2019TheRO, ; Rua2023ALE, ; Das2020CharacterizingTE, ; Das2016AQA, ). From these sources, we obtained a list of 87 keywords about Android performance issues. The keyword list includes terms like “slow response”, “memory leak”, “battery drain”, “cpu utilization”, “app size” and so on. The full list of keywords is available in (our_repo, ). These keywords were used to match the titles of user reviews, Stack Overflow posts, GitHub issues, and commits. It is important to note that title matching served only as an initial filtering step. In subsequent manual analysis, we further examined the associated descriptions and code snippets to extract actual performance issues and discover contributing factors, which may not have been explicitly mentioned in the original keyword list. This deeper analysis allowed us to go beyond the scope of the predefined keywords and uncover previously undocumented insights. Using these keywords, we filtered 60,684 user reviews to 165, 749,067 Stack Overflow discussions to 2,158, 16,977 GitHub issues down to 149, and 344,922 GitHub commits to 558.
@@
Ultimately, we obtained 114 user reviews, 1,484 Stack Overflow discussions, 69 GitHub issues, and 222 GitHub commits. Two of the authors performed the aforementioned manual checking individually, for any disagreement, they discussed their findings until reaching a consensus. For any persisting disagreement, a senior author made the final decision. We utilized Cohen’s kappa (Landis1977AnAO, ) to assess inter-rater agreement, yielding scores of 0.944 for user reviews, 0.868 for GitHub issues, 0.885 for GitHub commits, and 0.873 for Stack Overflow posts, all of which signify substantial agreement. We then manually identified the Android performance issues within this data and present our findings in the following section.
$\text{Google Play - User Reviews: 60,684}\rightarrow\text{165}$ $\rightarrow\text{114}$
$\text{Stack Overflow - Question Threads: 749,067}\rightarrow\text{2,158}$ $\rightarrow\text{1,484}$
$\text{GitHub - Issues \& Commits: 16,977/344,922}\rightarrow\text{149/558}$ $\rightarrow\text{69/222}$
@@ 2.2.1. Performance Issues in Real-world
Energy Consumption is also critical, as inefficient app power consumption can quickly drain batteries. These issues appeared in 20.5% of all reported reviews. E.g., one representative comment states: “severe battery drain on s10, causing 1% loss every 3 minutes. drain immediately disappeared after uninstall, back to 1% loss every 8-9 minutes.” This shows that performance issues can lead to users uninstalling the app. Storage Consumption is often complained by users that the size of the app is too large, or the cache space consumed is too large when using the app. These issues appeared in 9.0% of all reported reviews. As mobile devices have limited storage, excessive storage consumption can lead to problems like a lack of available space for other apps or files, and potential app crashes. E.g., one comment mentions: “..Every time a full size version of a Google app gets updated it eats up even more internal storage space… if there’s not enough storage space in your internal storage it won’t install certain updates or new apps from Google Play store, you have to uninstall updates or apps before reinstalling updates and apps…” Finally, Internet Data Usage issues are reported by users when apps consume excessive data, especially when they aren’t using Wi-Fi, making them more sensitive to data usage. For example, one user mentioned: “…Not sure why, but my data usage has skyrocketed. I’m on a limited plan, now I have to buy extra data every month…”
[Stack Overflow - Question Threads.] From the 1,484 question threads, we identified seven categories of performance issues, as illustrated in the Stack Overflow pie chart in Fig. 2. Interestingly, two performance issues (i.e., CPU usage and GPU usage) are not mentioned in user reviews. Additionally, while responsiveness is the most concerning issue for end users, it’s the second most frequent topic in developer discussions. This discrepancy indicates that the performance issues developers focus on and those users are concerned about do not align. This misalignment can be attributed to the differing perspectives of users and developers when evaluating performance. From a user-centric perspective, performance is often assessed through perceptible outcomes such as smooth responsiveness, battery consumption, storage consumption, and internet data usage. In contrast, a developer-centric perspective typically involves technical diagnostics, focusing on issues such as memory usage, CPU utilization, and GPU rendering efficiency. Since user satisfaction directly impacts app market share, developers should prioritize addressing the issues most important to users.
[Github - Issues & Commits.] The GitHub Issues and GitHub Commits pie charts in Fig. 2 show a similar distribution of performance issues across both data sources, with memory consumption, responsiveness, and storage consumption being the most frequently mentioned. Notably, while internet data usage is a niche issue mainly mentioned in user reviews and absent from GitHub commits, it also appears in GitHub issues, which offer valuable insights into the underlying causes discussed by developers.
@@
Table 1. Taxonomy of seven categories of performance issues and their contributing factors.
Category Factor Real-World Literature Problem Example
@@
Recursive Call Recursively sending requests rapidly creates excessive request queues and objects
@@
Image Decoding images without resizing; leaking image objects  (Song2021IMGDroidDI, ; Li2019CharacterizingAD, ; Carette2017InvestigatingTE, ; Kang2016ExperienceRD, ; Das2016AQA, ; MazueraRozo2020InvestigatingTA, )
@@
Resource Improper releasing of resources (e.g., Sensor, Camera)  (Liu2014GreenDroidAD, ; Liu2013WhereHM, ; Li2019DetectingAD, ; Behrouz2016EnergyawareTM, ; Li2022CombattingEI, ; Liu2016FixingRL, ; Couto2020EnergyRF, ; Jabbarvand2020AutomatedCO, ; Li2020EnergyDxDE, ; Lu2022DetectingRU, ; Banerjee2018EnergyPatchRR, ; Wu2016LightWeightIA, ; Jabbarvand2019SearchBasedET, ; Banerjee2014DetectingEB, ; Yan2013SystematicTF, ; Liu2019DroidLeaksAC, ; Das2016AQA, ; Wu2023ASL, )
@@
View Making allocations inside onDraw() routines  (Habchi2019TheRO, ; Couto2020EnergyRF, ; Habchi2020AndroidCS, )
@@
Data Type Declaring static variables retains strong references to previous tasks, preventing garbage collection (Das2020CharacterizingTE, ; Palomba2017LightweightDO, ; Habchi2020AndroidCS, )
@@
Inefficient Method Selecting a less efficient method introduces unnecessary delays in processing UI actions (Feng2019LearningPO, )
@@
Multi-Threads A thread gets blocked while waiting for a shared mutex held by another thread
@@
Resource Insufficient thread resource allocation causes slow response during simultaneous operations (Bhatt2020AutomatedRO, ; Prestat2024DynAMICSAT, )
Image Decoding large images in the UI thread (Song2021IMGDroidDI, ; Li2019CharacterizingAD, ; Carette2017InvestigatingTE, ; Kang2016ExperienceRD, ; Das2016AQA, )
@@
HTTP Request Overlong connection timeout keeps page in loading state (Zhao2018LeveragingPA, ; Das2016AQA, ; Zhao2019SystematicallyTA, ; Das2016AQA, )
UI Rendering Improper layout stacking causes redundant pixel redraws in the same frame (Gao2017EveryPC, ; Lin2024AgingOG, ; Liu2014CharacterizingAD, ; Habchi2019TheRO, ; Vsquez2015HowDD, ; Rua2023ALE, ; Das2020CharacterizingTE, ; Janssen2022OnTI, ; Nistor2014SunCatHD, ; MazueraRozo2020InvestigatingTA, ; Wu2023ASL, )
Loop An infinite while loop in the button’s onClick() event makes the app unresponsive after clicking (Lyu2018RemoveRF, ; Palomba2017LightweightDO, ; MazueraRozo2020InvestigatingTA, )
Main Thread Performing Heavy operations in the main thread  (Liu2014CharacterizingAD, ; Brocanelli2018HangDR, ; Lin2014RetrofittingCF, ; Gmez2016MiningTR, ; Vsquez2015HowDD, ; Das2020CharacterizingTE, ; Feng2019LearningPO, ; Das2016AQA, ; Wu2023ASL, )
Activity Loading new resources during activity restarts may freeze the UI (Guo2022DetectingAF, ; Chen2023TransparentRC, ; Lu2022DetectingRU, ; Callan2022ImprovingRO, )
AsyncTask Processing tasks in a single-threaded pool of AsyncTask since Android 3.0 (Lin2015StudyAR, ; Pan2020StaticAC, ; Kang2016DiagDroidAP, ; Fan2018EfficientlyMA, ; Prestat2024DynAMICSAT, ; Feng2019LearningPO, ; Das2016AQA, )
View Allocating new objects along with draw operations (Habchi2019TheRO, ; Couto2020EnergyRF, ; Das2020CharacterizingTE, ; Prestat2024DynAMICSAT, ; Wu2023ASL, )
@@
Message Backloging rapidly incoming messages in Message Handler’s single-threaded (Kang2016DiagDroidAP, )
Storage Consumption Cache Residual data remains after the application is uninstalled
Third-Party Library Unused third-party dependencies are packaged into the APK
Resource Unused resources (such as audio files and assets) are packaged into the APK
File Handling Temp files from URIs aren’t cleared after upload